OpenStudy (calculusxy):

Medal!!! A merchant places $1 in a jackpot on August 1, then draws the name of a regular customer. If the customer is present, the merchant adds $2 to the jackpot on August 2 and draws another name. Each day the merchant adds an amount equal to the day of the month. If the first person to win the jackpot wins $496, on what day of the month was her name drawn?

OpenStudy (calculusxy):

@hartnn

OpenStudy (calculusxy):

@rebeccaxhawaii

OpenStudy (goodfaze123):

hi

OpenStudy (calculusxy):

hi

OpenStudy (goodfaze123):

how you doing

OpenStudy (calculusxy):

fine.

OpenStudy (calculusxy):

can you help me with the question please?

OpenStudy (goodfaze123):

sure

OpenStudy (goodfaze123):

how many day are in august

OpenStudy (calculusxy):

31

OpenStudy (calculusxy):

It's okay. I got the answer: August 31

OpenStudy (mathmusician):

To quickly find the sum of a set of consecutive numbers, you can add the first and the last and multiply by half the number of values in the list. Based on that, we get the equation: (1 + x)(x/2) = 496 where x is the day of the award. x2 + x = 992 distribute and multiply both sides by 2 x2 + x - 992 = 0 (x + 32)(x - 31) = 0 factor and apply the zero product rule x = -32, 31 can't have negative days x = 31

OpenStudy (goodfaze123):

da fuk

OpenStudy (goodfaze123):

sorry a friend got got a hold of my keyboard